"A consumer has the utility function:u( x1, x2)=min(x1,x2)

and the endowment e=(2; 0).

Determine the consumer’s excess demand function!



1
Expert's answer
2021-03-17T08:32:13-0400

Total endowment"=" "2+0=2"

Excess supply"=x1-x2=x1"

Excess demand"=2(x1)=2x"